OpenStudy (anonymous):

Please help. Will give medal What equation in slope intercept form represents the line that passes through the two points? A. y = 3/7x -41/7 B. y = -7/3x -41/7 C. y = 7/3x +41/7 D. y = -3/7x + 41/7 (2, 5), (9, 2)

OpenStudy (anonymous):

so the first step is to work out the gradient (m) m = change in y / change in x so in this case, m = (5-2)/(2-9) m = (3)/(-7)

OpenStudy (anonymous):

Then, (y-b)=m(x-a) So substitute in your values (either point will work) (y-5)=(-3/7)(x-2) y - 5 = -3/7x +6/7 y = -3/7x + 41/7

OpenStudy (anonymous):

Thank you so much for helping me understand

OpenStudy (anonymous):

No problem :)